A truck of mass 50kg has its speed reduced from 4m/s to 1.5m/s in 30s. Find the braking force (assumed constant )
1
Expert's answer
2020-09-22T17:10:00-0400

"F=ma"


"v=v_0-at\to a=\frac{v_0-v}{t}=\frac{4-1.5}{30}=0.083(m/s^2)"


"F=50\cdot0.083=4.17(N)"